OpenStudy (anonymous):

integral cos x sin x sqrt(1-sin^4x) dx

OpenStudy (anonymous):

\[\int\limits \cos x \sin x \sqrt{1-\sin^4 x} dx\]

OpenStudy (anonymous):

really long and really ugly show steps http://www.wolframalpha.com/input/?i=cos+x+sin+x+sqrt%281-sin^4x%29+

OpenStudy (anonymous):

yeah, looks like it. is it the same for this one \[\int\limits (x \cos x + \sin x)\sqrt{1+x^2\sin^2x}\]